# Nonperturbative effects in $T\bar{T}$-deformed conformal field theories: A toy model for Planckian physics

### v1

@@ 3 Nonperturbative completion of two-point correlators
$\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\text{leading-log}}_{T\bar{T}}$ $=\frac{\Gamma(-\Delta+1)}{\pi 2^{2\Delta}\Gamma(\Delta)}\int_{-\infty}^{+\infty}d^{2}k|k|^{2(\Delta-1)}\exp^{i\vec{k}\cdot\vec{x}_{12}+\frac{\mu}{\pi}|k|^{2}\ln(|x_{12}|/\varepsilon)}\ ,$ (3.1)
@@
$\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\text{leading-log}}_{T\bar{T}}$ $=\sum_{n=0}^{\infty}(-1)^{n}\frac{(4\mu)^{n}}{n!\pi^{n}}\prod_{k=0}^{n-1}(\Delta+k)^{2}\frac{\ln^{n}(|x_{12}|/\varepsilon)}{|x_{12}|^{2(\Delta+n)}}\ .$ (3.2)
@@
Strictly speaking, the integral representation is convergent only in the range $10<{\rm Re}\Delta<$ 1. However, since the hypergeometric function itself is well-defined for generic $\Delta$, we analytically continue the result from this range to arbitrary values of $\Delta$, excluding $\Delta\in\mathbb{Z}_{+}$. For small $Z$, this expression admits a trans-series expansion:
$\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\rm leading{\text{-}}log}_{T{\bar{T}}}$ $={1\over|x_{12}|^{2\Delta}}\sum_{n=0}^{\infty}{(\Delta)_{n}(\Delta)_{n}\over n!}Z^{n}$ (3.14)
@@
$\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\rm leading{\text{-}}log}_{T{\bar{T}}}={(-1)^{\Delta}\over\pi 2^{2\Delta-1}\Gamma(\Delta)^{2}}\int_{-\infty}^{+\infty}d^{2}k|k|^{2(\Delta-1)}\ln|k|e^{i\vec{k}\cdot\vec{x}_{12}+{\mu\over\pi}|k|^{2}\ln(|x_{12}|/\varepsilon)}.$ (3.15)
@@
$\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\rm leading{\text{-}}log}_{T{\bar{T}}}$ $={(-1)^{\Delta}e^{-{1\over Z}}\over 4^{\Delta}\Gamma\left(\Delta\right)^{2}}\partial_{\Delta}\biggl{[}{\Gamma(\Delta)\over|x_{12}|^{2\Delta}}\left({4\over Z}\right)^{\Delta}\mbox{}_{1}F_{1}\left(1-\Delta;1;{1\over Z}\right)\biggr{]}$ (3.16)
@@
$\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\text{leading-log}}_{T{\bar{T}}}$ =$=\begin{cases}\dfrac{(-Z)^{-\Delta}}{|x_{12}|^{2\Delta}}U\left(\Delta,1,-\dfrac{1}{Z}\right)&\text{for }Z<0\\[4.30554pt] \dfrac{\Gamma(1-\Delta)Z^{-\Delta}}{|x_{12}|^{2\Delta}}e^{-{1\over Z}}\mbox{}_{1}F_{1}\left(1-\Delta;1;\dfrac{1}{Z}\right)&\text{for }Z>0\end{cases}$ (3.18)
@@ 4 Short-distance behavior and Planckian regimes
$\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\text{leading-log}}_{T{\bar{T}}}$$\sim\begin{cases}\dfrac{\Gamma(1-\Delta)}{\left(4\ell_{P}^{2}\ln(\ell_{P}/|x_{12}|)\right)^{\Delta}}&\text{for }\quad\mu>0\vspace{0.5em}\\ \dfrac{\left(\pi/4\ell_{P}^{2}\right)^{\Delta}}{\Gamma(\Delta)\left(\ln(\ell_{P}/|x_{12}|)\right)^{\Delta-1}}&\text{for }\quad\mu<0\ .\end{cases}$ (4.3)
@@ 5 Entanglement entropy
It is of interest to investigate how quantum entanglement responds to fluctuations in the underlying geometry, particularly when the entangling surfaces themselves are subject to fluctuation. [^5] For a discussion of the $T\bar{T}$ deformation of entanglement entropy on a cylinder, see [24]. Given the nonperturbative control of two-point functions, this question naturally presents itself as the next step to explore. As we will see, however, the results are somewhat puzzling: in one case, they broadly align with physical intuition, while in the other, they exhibit behavior that diverges markedly from what one might have anticipated. There is, however, an important caveat: our analysis is restricted to the leading logarithmic contribution, and subleading corrections might alter the overall picture.

### v2

@@ 3 Nonperturbative completion of two-point correlators
$\hskip-11.38092pt\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\text{leading-log}}_{T\bar{T}}$ $=\frac{\Gamma(-\Delta+1)}{\pi 2^{2\Delta}\Gamma(\Delta)}\int_{-\infty}^{+\infty}d^{2}k|k|^{2(\Delta-1)}\exp^{i\vec{k}\cdot\vec{x}_{12}+\frac{\mu}{\pi}|k|^{2}\ln(|x_{12}|/\varepsilon)}\ ,$ (3.1)
@@
$\hskip-11.38092pt\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\text{leading-log}}_{T\bar{T}}$ $=\sum_{n=0}^{\infty}(-1)^{n}\frac{(4\mu)^{n}}{n!\pi^{n}}\prod_{k=0}^{n-1}(\Delta+k)^{2}\frac{\ln^{n}(|x_{12}|/\varepsilon)}{|x_{12}|^{2(\Delta+n)}}\ .$ (3.2)
@@
Strictly speaking, the integral representation is convergent only in the range $0<{\rm Re}\Delta<1$. However, since the hypergeometric function itself is well-defined for generic $\Delta$, we analytically continue the result from this range to arbitrary values of $\Delta$, excluding $\Delta\in\mathbb{Z}_{+}$. For small $Z$, this expression admits a trans-series expansion:
$\hskip-14.22636pt\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\rm leading{\text{-}}log}_{T{\bar{T}}}$ $={1\over|x_{12}|^{2\Delta}}\sum_{n=0}^{\infty}{(\Delta)_{n}(\Delta)_{n}\over n!}Z^{n}$ (3.14)
@@
$\hskip-8.5359pt\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\rm leading{\text{-}}log}_{T{\bar{T}}}={(-1)^{\Delta}\over\pi 2^{2\Delta-1}\Gamma(\Delta)^{2}}\int_{-\infty}^{+\infty}d^{2}k|k|^{2(\Delta-1)}\ln|k|e^{i\vec{k}\cdot\vec{x}_{12}+{\mu\over\pi}|k|^{2}\ln(|x_{12}|/\varepsilon)}.$ (3.15)
@@
$\hskip 0.0pt\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\rm leading{\text{-}}log}_{T{\bar{T}}}$ $={(-1)^{\Delta}e^{-{1\over Z}}\over 4^{\Delta}\Gamma\left(\Delta\right)^{2}}\partial_{\Delta}\biggl{[}{\Gamma(\Delta)\over|x_{12}|^{2\Delta}}\left({4\over Z}\right)^{\Delta}\mbox{}_{1}F_{1}\left(1-\Delta;1;{1\over Z}\right)\biggr{]}$ (3.16)
@@
$\hskip-11.38092pt\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\text{leading-log}}_{T{\bar{T}}}$ =$=\begin{cases}\dfrac{(-Z)^{-\Delta}}{|x_{12}|^{2\Delta}}U\left(\Delta,1,-\dfrac{1}{Z}\right)&\text{for }Z<0\\[4.30554pt] \dfrac{\Gamma(1-\Delta)Z^{-\Delta}}{|x_{12}|^{2\Delta}}e^{-{1\over Z}}\mbox{}_{1}F_{1}\left(1-\Delta;1;\dfrac{1}{Z}\right)&\text{for }Z>0\end{cases}$ (3.18)
@@ 4 Short-distance behavior and Planckian regimes
$\langle{\cal O}_{\Delta}(x_{1}){\cal O}_{\Delta}(x_{2})\rangle^{\text{leading-log}}_{T{\bar{T}}}$$\sim\begin{cases}\dfrac{\Gamma(1-\Delta)}{\left(4\ell_{P}^{2}\ln(\ell_{P}/|x_{12}|)\right)^{\Delta}}&\text{for }\quad\mu>0\vskip 5.0pt\\ \dfrac{\left(\pi/4\ell_{P}^{2}\right)^{\Delta}}{\Gamma(\Delta)\left(\ln(\ell_{P}/|x_{12}|)\right)^{\Delta-1}}&\text{for }\quad\mu<0\ .\end{cases}$ (4.3)
@@ 5 Entanglement entropy
It is of interest to investigate how quantum entanglement responds to fluctuations in the underlying geometry, particularly when the entangling surfaces themselves are subject to fluctuation. [^5] For a discussion of the $T\bar{T}$ deformation of entanglement entropy on a cylinder, see [24]. Meanwhile, the $T\bar{T}$ deformation of Rényi entropy has been analyzed perturbatively to first order in [25, 26]. More recently, [27] investigates nonperturbative aspects of Rényi entropy in $T\bar{T}$-deformed CFTs, highlighting a potentially important role for instanton-like contributions. Given the nonperturbative control of two-point functions, this question naturally presents itself as the next step to explore. As we will see, however, the results are somewhat puzzling: in one case, they broadly align with physical intuition, while in the other, they exhibit behavior that diverges markedly from what one might have anticipated. There is, however, an important caveat: our analysis is restricted to the leading logarithmic contribution, and subleading corrections might alter the overall picture.